Suppose that LaTeX: X_{1}X 1 has a Poisson distribution with mean 2, LaTeX: X_{2}X 2has a Poisson distribution with mean 3 , LaTeX: X_{3}X 3 has a Poisson distribution with mean 5 and that LaTeX: X_{1}, X_{2} X 1 , X 2 and LaTeX: X_{3}X 3 are independent. Define LaTeX: Y = X_{1}+X_{2}+X_{3} Y = X 1 + X 2 + X 3. Determine the moment-generating function for Y.
Answer #1

Let , X1 → Poisson(mean = 11 = 2)

Therefore , the moment generating fucntion of X_1 is ,

My(t) = 42(t-1) = 2(t-1)

Also ,

Let , X2 + Poisson (mean = 12 = 3)

Therefore , the moment generating fucntion of X_2 is ,

Mx,(t) = e12(et-1) = 3(et-1

Let , X3 → Poisson (mean = 13 = 5)

Therefore , the moment generating fucntion of X_3 is ,

My(t) = e3(et-1) = €5(et-1)

Therefore , the moment generationg function of Y = X1 + X2 + X3 is ,

My(t) = MX1+Xz+xz(t)

= Mx,(t) * Mx (t) * Mx (t) ; Since , X_1 , X_2 and X_3 are independent.

= 2(t-1)* <3(et-1) * €5(et-1)

= (2+3+5)(e-1

= 10(-1

Therefore , Y → Poisson (1 = 10)